What quantity of 60% acid solution must be mixed with 30% acid solution to produce 300mL of a 50% solution?
Answer by Earlsdon(6294)

Let x = the number of ml of 60% acid solution to be added to (300-x) ml of 30% acid solution to obtain 300 ml of 50% acid solution. Change the percentages to their decimal equivalents.
0.6x + 0.3(300-x) = 0.5(300) Simplify and solve for x.
0.6x + 90 - 0.3x = 150
0.3x + 90 = 150
0.3x = 60
x = 200 ml